# Evaluating AI systems under uncertain ground truth: a case study in dermatology

### v1

@@ 2.4 Uncertainty-adjusted accuracy
Given a measure of annotation certainty, we intend to take it into account when evaluating performance of AI models. Specifically, we assume the label set $C$ used in Equation (5) to be a prediction set from a classifier – for example, corresponding to the top-$k$ logits of a deep neural networks. Then, we wish to measure the quality of this prediction set. If we knew the true plausibilities $\lambda^{*}=p(y|x)$, our ground truth target $Y_{\text{top}-j}(\lambda^{*})$ would be chosen as the top-$j$ elements of $\lambda^{*}$. The quality of the prediction can then be computed by evaluating the indicator of the event that the target set is contained within the prediction set, $\delta\left[{Y_{\text{top}-j}(\lambda^{*})\subseteq C_{\text{top}-k}(x)}\right]\in\{0,1\}$, assuming $j\leq k$ for simplicity). For example, the standard top-$k$ accuracy is an estimate of the probability where we specifically take $j=1$ and define
$\text{Accuracy}_{\text{top}-k}=p(Y_{\text{top}-1}(\lambda^{*})\subseteq C_{\text{top}-k}(x))=\mathbb{E}_{p(x)}\{\delta\left[{Y_{\text{top}-1}(\lambda^{*})\subseteq C_{\text{top}-k}(x)}\right]\}.$ (7)
However, we acknowledge that we cannot know $\lambda^{*}$ and that there is uncertainty on $\lambda$. Luckily, given $p(\lambda|b,x)$, we can quantify this uncertainty using Certainty from Equation (3). Integrating this into the above definition of accuracy yields the proposed uncertainty-adjusted version:
@@
Now, following (Wu and Crestani, 2003; Webber et al., 2010), we consider overlaps of increasingly larger prediction and ground truth sets and define uncertainty adjusted average overlap as
[…]
@@
Similar uncertainty-adjusted variants can be defined for other ranking-based metrics (Sakai, 2013), e.g., Kendall’s tau or Spearman’s footrule (Shieh, 1998; Fagin et al., 2003, 2004; Kumar and Vassilvitskii, 2010; Vigna, 2015). In general, we expect that uncertainty-adjusted variants can easily be defined for almost all relevant metrics in machine learning.
3 Case study: skin condition classification
Image and annotations
Refer to caption A0: {Pyogenic granuloma} {Hemangioma} {Melanoma} A1: {Angiokeratoma of skin} {Atypical Nevus} A2: {Hemangioma} {Melanocytic Nevus, Melanoma, O/E - ecchymoses present} A3: {Hemangioma, Melanoma, Skin Tag} A4: {Melanoma} A5: {Hemangioma} {Melanoma} {Melanocytic Nevus}
Plausibilities
Refer to captionRefer to caption
Pred.
Model A: $C_{\text{top-}3}=$Atyp. Nevus, Hemangioma, Melanocytic Nevus} Model B: $C_{\text{top-}3}=$Hemangioma, Melanocytic Nevus, Melanoma}
Eval.
Model A PL PrIRN $\mathbb{E}_{p(\lambda|b,x)}\{\delta[Y_{\text{top-}1}(\lambda)\subseteq C_{\text{top-}3}(x)]\}$ $0.7$ $0.52$ Model B PL PrIRN $\mathbb{E}_{p(\lambda|b,x)}\{\delta[Y_{\text{top-}1}(\lambda)\subseteq C_{\text{top-}3}(x)]\}$ $1$ $0.99$
Figure 5: An example from our case study: Given images (one shown), the task is to predict a set out of $419$ different skin conditions. For evaluation, six annotators provided (partial) rankings of potential ground truth conditions (top row). Given these annotations, our statistical aggregation models, PL and PrIRN, produce distributions over plausibilities (samples indicated in green), contrasted with deterministic aggregation such as IRN or ML (blue). These plausibilities represent categorical distributions over all conditions (second row). We then evaluate uncertainty-adjusted metrics based on these plausibilities: given prediction sets for models A and B (third row), we evaluate how often, in expectation, the top-1 label from the plausibilities is included (bottom row).
We demonstrate our framework on a case study in dermatology: skin condition classification from images (Liu et al., 2020). Here, several dermatologists provide annotations in the form of differential diagnoses, i.e., partial rankings of classes (see Figure 5). Based on previous work (Eng et al., 2019; Jain et al., 2021), we expect significant disagreement among annotators indicating a high degree of uncertainty in the ground truth. Following (Liu et al., 2020), we initially use inverse rank normalization (IRN) for deterministic aggregation. We then introduce two statistical aggregation models: a probabilistic version of IRN (PrIRN) as well as a Plackett–Luce (Plackett, 1975; Luce, 2012) (PL) based model. Both models include a reliability parameter described in Section 2 that reflects our trust in the annotators. Selecting a range of reliabilities, we evaluate annotation certainty as well as uncertainty-adjusted accuracy of classifiers in a range of different trust scenarios. Across reliabilities, we observe very high annotation uncertainty in the dataset of (Liu et al., 2020) and discuss implications for evaluating and selecting classifiers.
3.1 Dataset and methods
@@
The dataset of (Liu et al., 2020) includes $K=419$ different conditions which are to be predicted from three $448\times 448$ pixel images taken with consumer-grade cameras. This is illustrated in Figure 5 (top row), showing one of the input images alongside the corresponding annotations from six dermatologists. Each annotation includes a variable number of conditions combined with a confidence value. As these confidence values are not comparable across dermatologists, previous work (Eng et al., 2019; Roy et al., 2022; Azizi et al., 2022) uses them to obtain (partial) rankings of conditions (made explicit by braces in Figure 5). Details and statistics on the annotations are provided in Appendix C. Following (Liu et al., 2020), we use inverse rank normalization (IRN) to obtain a point estimate of plausibilities. Specifically, IRN weights each condition by its inverse rank, i.e., $\nicefrac{{1}}{{j}}$ for the $j$-th condition, sums across annotations and re-normalizes. We use these plausibilities to train several classifiers using the cross-entropy loss by varying architecture and hyper-parameters following (Roy et al., 2022). We randomly selected four of these classifiers for evaluation and comparison. In prior work, evaluation is based purely on the top-1 IRN label (i.e., the classifier is evaluated against the argmax of IRN plausibilities), ignoring both evidence of lower ranked conditions as well as the uncertainty induced through the aggregated annotations.
@@
Figure 6: For PL (top) and PrIRN (bottom), we plot annotation certainty as well as uncertainty-adjusted top-3 accuracy, top-2 and top-3 set accuracy and top-3 average overlap (y-axis). We consider four classifiers, A to D, across various reliabilities (x-axis). We omit exact reliability values as they are inherently not comparable between PL and PrIRN. Annotation certainty and evaluation metrics are averages across all examples and $M=1000$ plausibility samples. The shaded region additionally reports the standard deviation of metrics after averaging across plausibility samples; colored $\times$ mark performance against point estimates (ML for PL, IRN for PrIRN). We find that reliability severely impacts evaluated performance and that there is a significant variation induced by annotation uncertainty. Moreover, classifiers tend to perform significantly worse when evaluated against more than the top-1 labels using set accuracy.
In contrast, we propose two alternative statistical aggregation models following Section 2: a probabilistic version of IRN (PrIRN), and a Plackett–Luce (PL) based model. PrIRN interprets the IRN plausibilities $\hat{\lambda}$ as a maximum likelihood estimate of the parameters of a multinomial distribution. The posterior $p(\lambda|b)$ for sampling plausibilities is accordingly defined to be a Dirichlet distribution $\mathcal{D}(\gamma\hat{\lambda})$ where $\gamma$ is the aforementioned reliability parameter. Instead, PL is a standard model for rankings of conditions that we extend to handle partial rankings. PL assumes a categorical distribution over conditions from which annotators sample without replacement. Assuming independent gamma priors on plausibilities, a Gibbs sampler allows to efficiently sample plausibilities from the corresponding posterior. Similar to IRN, we consider the maximum likelihood (ML) estimate under PL likelihood as a deterministic counterpart to modeling the full posterior. Like PrIRN, PL model can also accommodate a reliability parameter that specifies our trust into the annotators. In both models, higher reliability will result in higher annotation certainty, cf. Equation (2). As it is generally unclear what the “right” reliability should be, we perform experiments across a range of reliabilities, corresponding to different scenarios of how much we trust our annotators. We refer to Appendix B for an in-depth description of both approaches, including several technical contributions in applying PL to partial rankings.
@@
Plausibilities from PrIRN and PL for a particular difficult case are shown in Figure 5 (second row). Compared to ML and IRN (blue bars), there is clearly significant variation in the sampled plausibilities (green dots) to the extent that the two most likely conditions (Hemangioma and Melanoma) may swap their positions. This also impacts evaluation: considering the two prediction sets from Figure 5 (third row), the first model (A) does not include both conditions, while the second model (B) does. Concretely, model A includes the top-1 label in only 70% of the sampled plausibilities, while model B always includes the top-1 label (100%), cf. Figure 5 (fourth row). In a nutshell, this summarizes our evaluation methodology on one specific case with fixed reliability, which we extend to the whole test set and across multiple reliabilities next.
3.2 Results
@@
Figure 7: For a fixed, medium reliability, we present annotation certainty and uncertainty-adjusted top-3 accuracy across examples and plausibility samples. Left: PL top-1 annotation certainty plotted against sorted examples (blue) in comparison to PrIRN (red). While there is high correlation between PL and PrIRN, there can be significant difference for individual examples. Middle: Uncertainty-adjusted top-3 accuracy for model A against sorted examples (blue). For many examples, the classifier does not consistently include all possible top-1 ground truth label in its predictions, resulting in values between 00 and $1$. Right: Histogram plot of uncertainty-adjusted top-3 accuracy (averaged over examples) across the $M=1000$ plausibility samples. Between worst- and best-case plausibilities, there can be up to 4% difference in accuracy.
@@
Our main results focus on evaluating (top-k) annotation certainty (see Equation (5)) alongside uncertainty-adjusted accuracy (see Equation (8)) across classifiers and reliabilities. As reliabilities are inherently not comparable across PL and PrIRN, we omit specific values. Then, Figure 6 (first column) highlights that average annotation certainty clearly increases with higher reliability; eventually being $1$ at infinite reliability. Also, top-$2$ and top-$3$ annotation certainty is significantly lower than top-1 annotation certainty, meaning that there is significant uncertainty not only in the top-1 condition but also lower ranked ones. This annotation uncertainty also has a clear impact on accuracy: Uncertainty-adjusted top-3 accuracy (second column) reduces significantly for lower reliability. As indicated by the shaded region, there is also high variation in accuracy across plausibility samples. In contrast, the ML and IRN plausibility point estimates, as used in previous work (Liu et al., 2020), typically overestimate performance and cannot provide an estimate of the expected variation in performance.
@@
In order to account for inherent uncertainty in the evaluation, i.e., plausibilities not being crisp distributions, we additionally consider set accuracy in Figure 6 (third and fourth columns). Strikingly, set accuracy is dramatically lower than standard accuracy, highlighting that the trained classifiers perform poorly on conditions likely ranked second or third. This is further emphasized by the fact that the reduction in accuracy is significantly larger than the corresponding drop in annotation certainty. Moreover, shifting focus to set accuracy also impacts the ranking of classifiers. This may has severe implications for hyper-parameter optimization and model selection which is typically based purely on standard accuracy. We also observe more significant differences between using PrIRN and PL, e.g., in terms of absolute accuracy numbers, their variation, or differences across classifiers. This highlights the impact that the statistical aggregation model can have on evaluation. As we might not want to put equal weight on all top-3 ground truth labels equally, Figure 6 (last column) also shows top-3 uncertainty-adjusted average overlap, where second and third condition are weighed by $\nicefrac{{1}}{{2}}$ and $\nicefrac{{1}}{{3}}$, respectively. Besides resulting in generally higher numbers, this also reduces variation significantly.
@@
Figure 8: Implications of ground truth uncertainty on risk categories. Right: Annotation certainty computed for risks in red in comparison to conditions blue. Going from 419 conditions to 3 risk levels clearly increases annotation certainty on average. But for many examples, risk categories remain very uncertain. Right: For many examples, there is evidence for multiple risk categories within the annotations. We plot a histogram of expected risk, the expected risk categories given the plausibilities, i.e., distributions over conditions.
@@
In the following, we focus on a fixed, medium reliability (as annotated in Figure 6) and consider annotation certainty and uncertainty-adjusted accuracy across examples. Specifically, Figure 7 (left) plots annotation certainty from PL (blue) and PrIRN (red) over examples (sorted for PL): For at least a quarter of the examples there is significant annotation uncertainty, i.e., top-1 annotation certainty is well below $1$. We also observe that annotation certainty is strongly correlated between PL and PrIRN (correlation coefficient 0.9). This indicates that similar examples are identified as having high annotation uncertainty. However, on individual examples, there can still be a significant difference. Similarly, we found that annotation certainty correlates well with annotator disagreement (see Appendix D). Figure 7 (middle) also shows uncertainty-adjusted top-3 accuracy against (sorted) examples. Again, for at least a quarter of examples, uncertainty-adjusted accuracy lies in between 00 and $1$, i.e., the top-3 prediction sets do not always include all possible top-1 ground truth labels. In Figure 7 (right), we also show results across plausibility samples might implicate in terms of aggregate statistics. While Figure 6 depicted the variability based on plausibility samples only through standard deviation error bars, these histograms clearly show that accuracy can easily vary by up to 4% between best and worst case.
@@
Besides performance evaluation across all 419 distinct conditions, previous work (Roy et al., 2022) also put significant focus on classifying risk categories, considering low, medium or high risk conditions. These categories are assigned to each condition independent of the actual case (e.g., Melanoma is a high-risk condition). As recommendations to users (e.g., whether the user should see a specialist) are similar for conditions in the same risk categories, it is often more important to correctly classify risk categories compared to individual conditions. Figure 8 (left), however, shows that these risk categories are also subject to significant uncertainty. This is made explicit by computing top-1 annotation certainty for risks categories (red) derived from the plausibilities over conditions. While this is generally higher than annotation certainty for conditions (blue) due to the smaller label space (3 risks vs. 419 conditions), annotation certainty remains low for many cases. This also has far-reaching consequences for evaluation. For example, evaluation metrics such as accuracy are often conditioned on high-risk cases. That is, for evaluation, we are interested in a classifier’s accuracy only considering high-risk cases. This conditioning, however, is not well defined in light of this uncertainty. This is made explicit in Figure 8 (right) which plots expected risk: the expected risk assignment for cases, based on plausibility samples, after mapping risk levels to an ordinal scale; low = 0, medium = 1, and high = 2. Most cases do not yield crisp risk assignments as there is typically evidence for multiple risk categories present in the annotations.

### v2

@@ 2.5 Uncertainty-adjusted accuracy
Given a measure of annotation certainty, we intend to take it into account when evaluating performance of AI models. Specifically, we assume the label set $C$ used in Equation (10) to be a prediction set from a classifier – for example, corresponding to the top-$k$ logits of a deep neural networks (denoted $C_{\text{top}-k}$. Then, we wish to measure the quality of this prediction set. If we knew the true plausibilities $\lambda^{*}=p(y|x)$, our ground truth target $Y_{\text{top}-j}(\lambda^{*})$ would be chosen as the top-$j$ elements of $\lambda^{*}$. The quality of the prediction can then be computed by evaluating the indicator of the event that the target set is contained within the prediction set, $\delta\left[{Y_{\text{top}-j}(\lambda^{*})\subseteq C_{\text{top}-k}(x)}\right]\in\{0,1\}$, assuming $j\leq k$ for simplicity). For example, the standard top-$k$ accuracy is an estimate of the probability where we specifically take $j=1$ and define
$\text{Accuracy}_{\text{top}-k}$ $=p(Y_{\text{top}-1}(\lambda^{*})\subseteq C_{\text{top}-k}(x))$ (12)
$=\mathbb{E}_{p(\lambda^{*},x)}\{\delta\left[{Y_{\text{top}-1}(\lambda^{*})\subseteq C_{\text{top}-k}(x)}\right]\}.$ (13)
However, we acknowledge that $\lambda^{*}$ is unknown and that there is uncertainty on $\lambda$. Given $p(\lambda|b,x)$, we can quantify this uncertainty using Certainty from Equation (8). Integrating this into the above definition of accuracy yields the proposed uncertainty-adjusted version:
@@
$=\mathbb{E}_{p(b,x)}\mathbb{E}_{p(\lambda|b,x)}\{\delta[Y_{\text{top-}1}(\lambda)\subseteq C_{\text{top-}k}(x)]\}.$ (15)
@@
$=\mathbb{E}_{p(b,x)}\mathbb{E}_{p(\lambda|b,x)}\{\delta\left[{Y_{\text{top}-|C|}(\lambda)=C(x)}\right]\}.$ (17)
@@
PL plausibilities                                                          PRIRN plausibilities
Refer to captionRefer to caption
Plausibilities
Pred.
Model A: $C_{\text{top-}3}=$
{Atyp. Nevus, Hemangioma, Melanocytic Nevus}
Model B: $C_{\text{top-}3}=$
{Hemangioma, Melanocytic Nevus, Melanoma}
Eval.
Model A PL PrIRN
$\mathbb{E}_{p(\lambda|b,x)}\{\delta[Y_{\text{top-}1}(\lambda)\subseteq C_{\text{top-}3}(x)]\}$ $0.7$ $0.52$
Model B PL PrIRN
$\mathbb{E}_{p(\lambda|b,x)}\{\delta[Y_{\text{top-}1}(\lambda)\subseteq C_{\text{top-}3}(x)]\}$ $1$ $0.99$
Fig. 5: Plausibilities and evaluation results for the example and annotations from Figure 3. Our statistical aggregation models, PL and PrIRN, produce distributions over plausibilities (samples indicated in green, medium reliability), contrasted with deterministic aggregation such as IRN or ML (blue). These plausibilities represent categorical distributions over all conditions (first row). We then evaluate uncertainty-adjusted metrics based on these plausibilities: given prediction sets for models A and B (second row), we evaluate how often, in expectation, the top-1 label from the plausibilities is included (third row).
@@
Following [114, 111], we also consider overlaps of increasingly larger prediction and ground truth sets and define uncertainty adjusted average overlap as
$\begin{split}\text{UA-AverageOverlap@L}\hskip 85.35826pt\\ =\mathbb{E}_{p(b,x)}\Big{\{}\tfrac{1}{L}\sum_{k=1}^{L}\text{UA-Overlap}(C_{\text{top-}k}(x),x)\Big{\}}.\end{split}$ (20)
@@
Note that using this formulation also avoids difficulties computing average overlap directly against partial rankings, see C. Similar uncertainty-adjusted variants can be defined for other ranking-based metrics [86], e.g., Kendall’s tau or Spearman’s footrule [96, 32, 33, 53, 106]. In general, we expect that uncertainty-adjusted variants can easily be defined for almost all relevant metrics in machine learning.
3 Experiments
We present comprehensive experiments on the skin condition classification data of [58], where we expect high annotation uncertainty across a significant proportion of the dataset. As outlined in Section 2.3, we initially use inverse rank normalization (IRN) for deterministic aggregation. We experiment with our two statistical aggregation models, probabilistic IRN (PrIRN) and Plackett–Luce [75, 61] (PL). Both models include a reliability parameter, described in Section 2, that reflects our trust in the annotators. While domain expertise might help select an appropriate rater reliability, we consider the case where it may be unclear what the “right” reliability for our statistical aggregation models should be. Thus, we conduct experiments across a range of reliabilities, exploring different levels of annotator trust.
3.1 Dataset, Model and Example
@@
The dataset of [58] includes $K=419$ different conditions which are to be predicted from three $448\times 448$ pixel images taken with consumer-grade cameras. Each annotation includes a variable number of conditions combined with a confidence value. As these confidence values are not comparable across dermatologists, previous work [30, 84, 5] uses them to obtain differential diagnoses, formalized as partial rankings of conditions. Details and statistics on the annotations are provided in D. We train several classifiers using cross-entropy against the IRN plausibilities by randomly varying architecture and hyper-parameters and picking four models for our experiments. We follow the exact same set up as described in detail in [84] and D.2. We randomly selected four of these classifiers for evaluation and comparison.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Fig. 6: For PL (top) and PrIRN (bottom), we plot annotation certainty as well as uncertainty-adjusted top-3 accuracy, top-2 and top-3 set accuracy and top-3 average overlap (y-axis). We consider four classifiers, A to D, across various reliabilities (x-axis). We omit exact reliability values as they are inherently not comparable between PL and PrIRN. Annotation certainty and evaluation metrics are averages across all examples and $M=1000$ plausibility samples. The shaded region additionally reports the standard deviation of metrics after averaging across plausibility samples; colored $\times$ mark performance against point estimates (ML for PL, IRN for PrIRN). We find that reliability severely impacts evaluated performance and that there is a significant variation induced by annotation uncertainty. Classifiers tend to perform significantly worse when evaluated against more than the top-1 labels using set accuracy.
@@
Plausibilities from PrIRN and PL for the case of Figure 3 are shown in Figure 5 (top row). Compared to ML and IRN (blue bars), there is clearly significant variation in the sampled plausibilities (green dots) to the extent that the two most likely conditions (Hemangioma and Melanoma) may swap their positions. Note that some plausibility samples can change the order of conditions radically from their deterministic, i.e., infinite reliability, counterparts (ML for PL and IRN for PrIRN). The example also highlights differences between PL and PrIRN. As IRN resolves ties deterministically by equally distributing weight across tied conditions, the rank of conditions such as “Melanocytic Nevus” that are involved in ties can be unintuitive after aggregation. This also impacts evaluation: considering the two prediction sets from Figure 5 (second row), the first model (A) does not include both conditions, while the second model (B) does. Concretely, model A includes the top-1 label in only 70% of the sampled plausibilities, while model B always includes the top-1 label (100%), cf. Figure 5 (third row). In a nutshell, this summarizes our evaluation methodology on one specific case with fixed reliability, which we extend to the whole test set and across multiple reliabilities next.
3.2 Main results
@@
Fig. 7: For a fixed, medium reliability, we present annotation certainty and uncertainty-adjusted top-3 accuracy across examples and plausibility samples. Left: PL top-1 annotation certainty plotted against sorted examples (blue) in comparison to PrIRN (red). While there is high correlation between PL and PrIRN, there can be significant difference for individual examples. Middle: Uncertainty-adjusted top-3 accuracy for model A against sorted examples (blue). For many examples, the classifier does not consistently include all possible top-1 ground truth label in its predictions, resulting in values between 00 and $1$. Note that top-1 certainties are sorted separately. Right: Histogram plot of uncertainty-adjusted top-3 accuracy (averaged over examples) across the $M=1000$ plausibility samples. Between worst- and best-case plausibilities, there can be up to 4% difference in accuracy.
@@
Our main results focus on evaluating (top-k) annotation certainty alongside uncertainty-adjusted accuracy across classifiers and reliabilities. As reliabilities are inherently not comparable across PL and PrIRN, we omit specific values. Then, Figure 6 (first column) highlights that average annotation certainty clearly increases with higher reliability; eventually being $1$ at infinite reliability. Also, top-$2$ and top-$3$ annotation certainty is significantly lower than top-$1$ annotation certainty, meaning that there is significant uncertainty not only in the top-$1$ condition but also lower ranked ones. This annotation uncertainty also has a clear impact on accuracy: Uncertainty-adjusted top-$3$ accuracy (second column) reduces significantly for lower reliability. As indicated by the shaded region, there is also high variation in accuracy across plausibility samples. In contrast, the ML and IRN plausibility point estimates, as used in previous work [58], typically overestimate performance and cannot provide an estimate of the expected variation in performance.
@@
In order to account for inherent uncertainty in the evaluation, we additionally consider set accuracy in Figure 6 (third and fourth columns). Strikingly, set accuracy is dramatically lower than standard accuracy, highlighting that the trained classifiers perform poorly on conditions likely ranked second or third. This is further emphasized by the fact that the reduction in accuracy is significantly larger than the corresponding drop in annotation certainty. Moreover, shifting focus to set accuracy also impacts the ranking of classifiers. This can have severe implications for hyper-parameter optimization and model selection which is typically based purely on standard accuracy at infinite reliability. We also observe more significant differences between using PrIRN and PL, e.g., in terms of absolute accuracy numbers, their variation, or differences across classifiers. This highlights the impact that the statistical aggregation model can have on evaluation. Lastly, Figure 6 (last column) also shows top-3 uncertainty-adjusted average overlap, where second and third conditions are weighed by $\nicefrac{{1}}{{2}}$ and $\nicefrac{{1}}{{3}}$, respectively. Besides resulting in generally higher numbers, this also reduces variation significantly.
PL risk and condition annotation certainty                        Expected risk from PL plausibilities
@@
Fig. 8: Implications of ground truth uncertainty on risk categories. Right: Annotation certainty computed for risks in red in comparison to conditions blue. Going from 419 conditions to 3 risk levels clearly increases annotation certainty on average. But for many examples, risk categories remain very uncertain. Right: For many examples, there is evidence for multiple risk categories within the annotations. We plot a histogram of expected risk, the mean/max/min expected risk category given the plausibilities, i.e., distributions over conditions, after mapping risk levels to low = 0, medium = 1 and high = 2 (see text).
3.3 Analysis
@@
In the following, we focus on a fixed, medium reliability (as annotated in Figure 6) and consider annotation certainty and uncertainty-adjusted accuracy across examples. Specifically, Figure 7 (left) plots annotation certainty from PL (blue) and PrIRN (red) over examples (sorted for PL): For at least a quarter of the examples there is significant annotation uncertainty, i.e., top-$1$ annotation certainty is well below $1$. We also observe that annotation certainty is strongly correlated between PL and PrIRN (correlation coefficient 0.9). This indicates that similar examples are identified as having high annotation uncertainty. However, on individual examples, there can still be a significant difference. Similarly, we found that annotation certainty correlates well with annotator disagreement (see E). Figure 7 (middle) also shows uncertainty-adjusted top-3 accuracy against (sorted) examples. Again, for at least a quarter of examples, uncertainty-adjusted accuracy is well below $1$, i.e., the top-$3$ prediction sets do not always include all possible top-1 ground truth labels. In Figure 7 (right), we also show results across plausibility samples. While Figure 6 depicted the variability based on plausibility samples only through standard deviation error bars, these histograms clearly show that accuracy can easily vary by up to 4% between best and worst case.
@@
Besides performance evaluation across all $419$ distinct conditions, previous work [84] also put significant focus on classifying risk categories, considering low, medium or high risk conditions. These categories are assigned to each condition independent of the actual case (e.g., Melanoma is a high-risk condition). As recommendations to users (e.g., whether the user should see a specialist) are similar for conditions in the same risk categories, it is often more important to correctly classify risk categories compared to individual conditions. Figure 8 (left), however, shows that these risk categories are also subject to significant uncertainty. This is made explicit by computing top-$1$ annotation certainty for risk categories (red) derived from the plausibilities over conditions. While this is generally higher than annotation certainty for conditions (blue) due to the smaller label space (3 risks vs. 419 conditions), annotation certainty remains low for many cases. This also has far-reaching consequences for evaluation. For example, evaluation metrics such as accuracy are often conditioned on high-risk cases. That is, for evaluation, we are interested in a classifier’s accuracy only considering high-risk cases. This conditioning, however, is not well defined in light of this uncertainty. This is made explicit in Figure 8 (right) which plots expected risk: the expected risk assignment for cases, based on plausibility samples, after mapping risk levels to an ordinal scale; low = 0, medium = 1, and high = 2. That is, taking the expectation of the risk level (0, 1, or 2) over a plausibility sample. The figure then shows mean/max/min across our plausibility samples. Most cases do not yield crisp risk assignments as there is typically evidence for multiple risk categories present in the annotations.